Parth (parthkohli):

How do you derive this?\[ \sum_{i = 1}^{n} i^2 = {n(n + 1)(2n + 1) \over 6}\]

Parth (parthkohli):

Hello Sir Elias, can you explain it to me in simple terms?

OpenStudy (anonymous):

By induction. I think that I did this before on this site.

Parth (parthkohli):

Is there no other direct proof?

OpenStudy (anonymous):

I do not know another one.

Parth (parthkohli):

OK. Can you help me prove it? I did it already for 1. Now for the difficult parts: \(k\) and \(k + 1\).

OpenStudy (anonymous):

\[(n+1)^3-n^3=3n^2+3n+1\]\[(n)^3-(n-1)^3=3(n-1)^2+3(n-1)+1\]\[(n-1)^3-(n-2)^3=3(n-2)^2+3(n-2)+1\]... ... \[(2)^3-(1)^3=3(1)^2+3(1)+1\]add them all LHS's and RHS's

Parth (parthkohli):

... @mukushla Thank you. Can you please explain?

OpenStudy (anonymous):

we know that\[(n+1)^3-n^3=3n^2+3n+1\]

Parth (parthkohli):

But, why are we doing \((n + 1)^3\) when that is completely out of subject?

OpenStudy (anonymous):

when we add LHS's they will cancel each other

OpenStudy (anonymous):

and because that difference of cubes generates n^2 terms for us

Parth (parthkohli):

I still don't get it :( Perhaps because I'm not a genius like you guys =D

OpenStudy (anonymous):

no thats because i cant explain well :(

OpenStudy (amistre64):

0 0 [1] 5 14 30 55 91 ; squares added 0 1 [4] 9 16 25 36 ; odds added euqls squares 1 3 [5] 7 9 11 ; odds 2 2 [2] 2 2 ; constants of 2s this setup is related to pascals triangle, and when a sequence has multiple tiers of differences, till it reaches a constant, we can form a rule for it using those [..] as coeefs \[a\frac{x^0}{0!}+b\frac{n}{1!}+c\frac{n(n-1)}{2!}+d\frac{n(n-1)(n-2)}{3!}+...\] \[a=1;~b=4;~c=5;~d=2\]

OpenStudy (amistre64):

ignore the x^0, its left over froma bad idea lol

OpenStudy (amistre64):

if we turn these rows of differences clockwise 90 degrees we get the relation to the pascals triangle 1 2 1 2 3 1 2 5 4 1 2 7 9 5 1 i spose this should be flipped about so the 1s are on the other side to be consistent with the coeffs

OpenStudy (amistre64):

and to note, my setup starts with n=0; but can be modified for n=1 by subtracting 1 from all the n parts to adjust

OpenStudy (amistre64):

or just use the 5,9,7,2 instead :)

OpenStudy (amistre64):

ugh, went the wrong way ... 0,1,3,2 for n=1,2,3,... http://www.wolframalpha.com/input/?i=n%2B3n%28n-1%29%2F2%2B2n%28n-1%29%28n-2%29%2F6

OpenStudy (experimentx):

you forgot this picture http://jeremykun.files.wordpress.com/2011/06/triangle-proof.png

Parth (parthkohli):

Oh, yeah. lol